RAITH status Monday March 1, 2004: 17:21 hours.
jwc at snf.stanford.edu
Mon Mar 1 17:32:54 PST 2004
Neither Mike Santomango nor I were pleased with the emission and beam
profiles of the FE-GUN Tip as we arrived in this morning. While the
emission is near normal levels we had to drive the FE-gun far over
specifications to achieve the normal level. EHT V at 6400 V instead of
the normal 4400 - 5400 V.
We have decided the best approach at this point is to simply replace the
FE-GUN tip with a new clean unused tip and perform another bake out and
run up sequence. While we were open we also cleaned the Extractor plate
of Hydrocarbon residue that appeared over the weekend.
We just started pumping and will perform an extended bake out until
WEDNESDAY morning. Wednesday afternoon we will perform alignments and
adjustments and will again attempt to qualify the systems for line width
and stitching performance. I would expect access to reopen to users
Thursday morning. Please adjust your plans accordingly.
More information about the raith